# Gauges and Accelerated Optimization over Smooth and/or Strongly Convex Sets

### v3

@@ A.2.3 Proof of Proposition 3
$\mathbf{(b)\Rightarrow(a)}$. We prove the contrapositive. Suppose $B=B(x-\frac{1}{\beta}\zeta_{x},\frac{1}{\beta})\not\subseteq S$. Then the nearest boundary point to $x-\frac{1}{\beta}\zeta_{x}$, denoted $y=\operatornamewithlimits{argmin}_{v\in\mathrm{bdry\ }S}{\|v-(x-\frac{1}{\beta}\zeta_{x})\|}$, must have $\|y-(x-\frac{1}{\beta}\zeta_{x})\|<1/\beta$. The first-order optimality then ensures $\zeta_{y}=\frac{y-(x-\frac{1}{\beta}\zeta_{x})}{\|y-(x-\frac{1}{\beta}\zeta_{x})\|}\in N_{S}(y)$. Let $z=x-\frac{1}{\beta}\zeta_{x}+\frac{1}{\beta}\zeta_{y}$. Since $z-y$ is a normal vector of $S$ at $y$, one has $(z-y)^{T}(y-x)\geq 0$. Then noting $\|z-y\|>0$, we have
$\|z-x\|^{2}=\|z-y\|^{2}+2(z-y)^{T}(y-x)+\|y-x\|^{2}-2>\|y-x\|^{2}\ .$
Hence $\|y-x\|<\|z-x\|=\frac{1}{\beta}\|\zeta_{x}-\zeta_{y}\|$. Then Cauchy-Schwartz implies $\|\zeta_{x}-\zeta_{y}\|^{2}>\beta(\zeta_{x}-\zeta_{y})^{T}(x-y)$.
$\mathbf{(a)\Rightarrow(c)}$. Define the Minkowski Difference of two sets as $A-B=\{x\in\mathcal{E}:x+B\subseteq A\}$ and the support function of a set as $\sigma_{S}(a)=\sup_{x\in S}a^{T}x$. Note for closed convex sets $A,B$, the support function has
@@
Letting $S_{0}=S-B(0,\frac{1}{\beta})$, it is immediate that $S_{0}+B(0,\frac{1}{\beta})\subseteq S$. Then we need to show $S_{0}+B(0,\frac{1}{\beta})\supseteq S$. By properties (39) and (40), it suffices to show $\sigma_{S_{0}}(d)+\sigma_{B(0,\frac{1}{\beta})}(d)\geq\sigma_{S}(d)$ for every $d\in\mathcal{E}$.
Suppose $\sigma_{S}(d)$ is finite. Then, there exists a sequence $\{x_{i}\}$ with $x_{i}\in\mathrm{bdry\ }S$ such that $\lim_{i\rightarrow\infty}d^{T}x_{i}=\sigma_{S}(d)$. For each $x_{i}$, letting $\zeta_{i}\in N_{S}(x_{i})$ be some unit normal vector, the assumed smoothness ensures $B(x_{i}-\frac{1}{\beta}\zeta_{i},\frac{1}{\beta})\subseteq S$. Hence $x_{i}-\frac{1}{\beta}\zeta_{i}\in S-B(0,\frac{1}{\beta})=S_{0}$. So
@@
Next, suppose $\sigma_{S}(d)=\infty$. Let $z\in\mathrm{bdry\ }S$ be some point on the boundary of $S$. Then, there exists a ray $\{z+t\Delta\mid t\geq 0\}\subseteq S$ with $d^{T}\Delta>0$. Letting $\zeta_{z}\in N_{S}(z)$ be a unit normal vector, we have $B(z-\frac{1}{\beta}\zeta_{z},\frac{1}{\beta})\subseteq S$. By convexity, one has $\cup_{t\geq 0}B(z-\frac{1}{\beta}\zeta_{z}+t\Delta,\frac{1}{\beta})\subseteq S$. Consequently, $z-\frac{1}{\beta}\zeta_{z}+t\Delta\in S_{0}$ for any $t\geq 0$. Thus
@@
A.2.4 Proof of Proposition 4
The equivalences of $\mathbf{(a)},\mathbf{(b)},\mathbf{(c)},\mathbf{(d)}$ are given by the equivalences of $\mathbf{(e),(h),(f),(l)}$ in [25, Theorem 2.1, page 265-266].
A.2.5 Proof of Lemma 7
This result follows rather directly from relating the local quadratic upper and lower bounds from smoothness and strong convexity of $h$ to balls inside its epigraph. Note given any $g\in\partial h(x)$, $(\zeta,\delta)=(g,-1)/\|(g,-1)\|$ is a unit normal $(\zeta,\delta)\in N_{\mathrm{epi\ }h}(x,h(x))$. First, we address smoothness. Consider any $\tilde{L}>L$ and $\eta>0$ such that all $y\in B(x,\eta)$ have
@@ A.2.5 Proof of Lemma 7
Local $\tilde{\beta}$-smoothness of epi $\text{epi }h$ w.r.t. $((x,h(x)),(\zeta,\delta))$ corresponds to the following upper bound holding for all $y$ near $x$
@@
$\nabla^{2}q(x)$ $=\tilde{L}I\qquad\text{and}\qquad\nabla^{2}b(x)=\frac{\tilde{\beta}}{|\delta|}I+\frac{\tilde{\beta}\zeta\zeta^{T}}{|\delta|^{3}}\ .$
Thus $\nabla^{2}q(x)-\nabla^{2}b(x)$ is positive semidefinite whenever $\tilde{L}-\tilde{\beta}/|\delta|\geq 0$. So local smoothness holds with any $\tilde{\beta}\geq\tilde{L}\delta$-smoothness holds, and so $L\delta$ local smoothness holds.
@@
Noting $\|\zeta\|^{2}=1-\delta^{2}$, the minimum eigenvalue of $\nabla^{2}q(x)-\nabla^{2}b(x)$ is $\tilde{\mu}-\tilde{\alpha}/\delta^{3}$ and so any $\tilde{\alpha}\leq\tilde{\mu}\delta^{3}$-strong convexity of epi $\text{epi }h$ w.r.t. $((x,h(x)),(\zeta,\delta))$ holds locally.
A.2.6 Proof of Lemma 19
Consider any $y\in\mathbb{E}$ with $\gamma_{S}(y)\neq 0$. Let $\bar{y}=y/\gamma_{S}(y)$ be the corresponding point on the boundary of $S$. Note that $\zeta\in N_{S}(\bar{y})$ if and only if the halfspace =$\mathcal{H}=\{z\mid\zeta^{T}(z-\bar{y})\leq 0\}\supseteq{S}$, or equivalently $\gamma_{S}\geq\gamma_{\mathcal{H}}$. Then it yields
$N_{S}(\bar{y})$ =$=\{\zeta\mid\gamma_{S}(z)\geq\gamma_{\mathcal{H}}(z),\forall z\}$
=$=\{\zeta\mid\gamma_{S}(z)\geq\gamma_{S}(\bar{y})+\frac{\zeta^{T}(z-\bar{y})}{\zeta^{T}\bar{y}},\forall z\}$
=$=\{\zeta\mid\gamma_{S}(z)\geq\gamma_{S}(y)+\frac{\zeta^{T}(z-y)}{\zeta^{T}\bar{y}},\forall z\}$
=$=\{\zeta\mid\frac{\zeta}{\zeta^{T}\bar{y}}\in\partial\gamma_{S}(y)\}$
where the second equality is given by Example 3.1, $\gamma_{\mathcal{H}}(z)=\frac{\zeta^{T}z}{\zeta^{T}\bar{y}}=\gamma_{S}(\bar{y})+\frac{\zeta^{T}(z-\bar{y})}{\zeta^{T}\bar{y}}$, and the third equality is by rescaling. Then by the chain rule and this gauge subgradient formula,
@@ A.2.6 Proof of Lemma 19
Then the gauge of this ball is given by
@@
$=\inf\{\lambda>0\mid(\bar{2}\zeta^{T}\bar{y}-\|\bar{y}\|^{2})\lambda^{2}+2\lambda(\bar{y}-\bar{\zeta})^{T}z-\|z\|^{2}\geq 0\}$
[…]
@@
    $\nabla^{2}\frac{1}{2}\gamma_{B}^{2}(y)=\frac{1}{(\bar{\zeta}^{T}\bar{y})^{2}}\left(3\bar{\zeta}\bar{\zeta}^{T}-(\bar{\zeta}\bar{y}^{T}+\bar{y}\bar{\zeta}^{T})+\bar{\zeta}\bar{y}\right)\ .$
Case 2: $\|\bar{y}\|^{2}\neq 2\zeta^{T}\bar{y}$
@@
    $\quad\left.-\frac{\left(2\left((\bar{y}-\bar{\zeta})^{T}z\right)^{2}+(2\bar{\zeta}^{T}\bar{y}-\|\bar{y}\|^{2})\|z\|^{2}\right)(\bar{y}-\bar{\zeta})+(2\bar{\zeta}^{T}\bar{y}-\|\bar{y}\|^{2})(\bar{y}-\bar{\zeta})^{T}zz}{\sqrt{\left((\bar{y}-\bar{\zeta})^{T}z\right)^{2}+(2\bar{\zeta}^{T}\bar{y}-\|\bar{y}\|^{2})\|z\|^{2}}}\right)\ ,$
@@
    $\quad-\frac{\|z\|^{2}\left((\bar{y}-\bar{\zeta})z^{T}+z(\bar{y}-\bar{\zeta})^{T}\right)}{\left(\left((\bar{y}-\bar{\zeta})^{T}z\right)^{2}+(2\bar{\zeta}^{T}\bar{y}-\|\bar{y}\|^{2})\|z\|^{2}\right)^{\frac{3}{2}}}+\frac{(\bar{y}-\bar{\zeta})^{T}zzz^{T}}{\left(\left((\bar{y}-\bar{\zeta})^{T}z\right)^{2}+(2\bar{\zeta}^{T}\bar{y}-\|\bar{y}\|^{2})\|z\|^{2}\right)^{\frac{3}{2}}}$
    $\quad+\left(\frac{1}{2\bar{\zeta}^{T}\bar{y}-\|\bar{y}\|^{2}}-\frac{(\bar{y}-\bar{\zeta})^{T}z}{(2\bar{\zeta}^{T}\bar{y}-\|\bar{y}\|^{2})\sqrt{\left((\bar{y}-\bar{\zeta})^{T}z\right)^{2}+(2\bar{\zeta}^{T}\bar{y}-\|\bar{y}\|^{2})\|z\|^{2}}}\right)I\ .$
    At point $z=y$, the numerator in the above Hessian formula can be notably simplified since $((\bar{y}-\bar{\zeta})^{T}y)^{2}+(2\bar{\zeta}^{T}\bar{y}-\|\bar{y}\|^{2})\|y\|^{2}=(\bar{\zeta}^{T}\bar{y})^{2}$. After such simplifications, the Hessian matrix at $y$ is given by
@@ A.2.7 Proof of Lemma 20
$\lambda_{i}$ $=0,\qquad\text{for }i=2,...,n-1$
@@
$\lambda^{2}-\left(C_{1}\|a\|^{2}+2C_{2}a^{T}b+C_{3}\|b\|^{2}\right)\lambda+\left(C_{1}C_{3}-C_{2}^{2}\right)\left(\|a\|^{2}\|b\|^{2}-(a^{T}b)^{2})\right)=0\ .$
If $a$ and $b$ are linearly dependent, then the matrix is a rank one matrix. The only nonzero eigenvalue is $\lambda=C_{1}\|a\|^{2}+2C_{2}a^{T}b+C_{3}\|b\|^{2}$, which matches the formula with $\lambda_{1}=\lambda_{2}=...=\lambda_{n-1}=0$. ∎
@@
$\lambda_{i}=$ $0+\frac{1}{\bar{\zeta}^{T}\bar{y}}=\frac{1}{\bar{\zeta}^{T}\bar{y}},\qquad\text{for }i=2,...,n-1\ .$

### v4

@@ A.2.3 Proof of Proposition 3
$\mathbf{(b)\Rightarrow(a)}$. Fix $x\in\mathrm{bdry\ }S$ and a unit $\zeta_{x}\in N_{S}(x)$, and set $c:=x-\frac{1}{\beta}\zeta_{x}$. We prove $B(c,1/\beta)\subseteq S$ by contradiction. Supposing not, then there exists $y\in\mathrm{bdry\ }S$ with $d:=\|c-y\|<1/\beta$. Choose $y$ to be a nearest boundary point to $c$ (i.e., $y\in\arg\min_{v\in\mathrm{bdry\ }S}\|v-c\|$). By optimality and convexity of $S$,
$(z-y)^{T}(c-y)\leq 0\quad\forall z\in S\qquad\Rightarrow\qquad\zeta_{y}:=\frac{c-y}{\|c-y\|}\in N_{S}(y),$
and $\zeta_{y}$ is unit. Apply (b) to $x,y$ and $\zeta_{x},\zeta_{y}$:
$2(1-\zeta_{x}^{T}\zeta_{y})\ \leq\ \beta\big(\zeta_{x}-\zeta_{y}\big)^{T}(x-y)=\beta\Big(\tfrac{1}{\beta}+\zeta_{x}^{T}(c-y)-\tfrac{1}{\beta}\,\zeta_{x}^{T}\zeta_{y}-d\Big)=(1-\zeta_{x}^{T}\zeta_{y})+\beta\big(\zeta_{x}^{T}(c-y)-d\big).$
Hence
$1-\zeta_{x}^{T}\zeta_{y}\leq\beta\big(\zeta_{x}^{T}(c-y)-d\big).$ (45)
By Cauchy–Schwarz, $\zeta_{x}^{T}(c-y)\leq\|c-y\|=d$, so the right-hand side of (45) is $\leq 0$. Therefore $1-\zeta_{x}^{T}\zeta_{y}\leq 0$, i.e., $\zeta_{y}=\zeta_{x}$, and equality must hold in Cauchy–Schwarz, so $\zeta_{x}^{T}(c-y)=d$.
But $\zeta_{y}=\zeta_{x}\in N_{S}(x)\cap N_{S}(y)$ implies $\zeta_{x}^{T}x=\sigma_{S}(\zeta_{x})=\zeta_{x}^{T}y$. Thus
$\zeta_{x}^{T}(c-y)=\zeta_{x}^{T}(x-\tfrac{1}{\beta}\zeta_{x}-y)=\underbrace{\zeta_{x}^{T}x-\zeta_{x}^{T}y}_{=\,0}-\tfrac{1}{\beta}=-\tfrac{1}{\beta},$
which contradicts $\zeta_{x}^{T}(c-y)=d>0$.
$\mathbf{(a)\Rightarrow(c)}$. Define the Minkowski Difference of two sets as $A\ominus B=\{x\in\mathcal{E}:x+B\subseteq A\}$ and the support function of a set as $\sigma_{S}(a)=\sup_{x\in S}a^{T}x$. Note for closed convex sets $A,B$, the support function has
@@
Letting $S_{0}=S\ominus B(0,\frac{1}{\beta})$, it is immediate that $S_{0}+B(0,\frac{1}{\beta})\subseteq S$. Then we need to show $S_{0}+B(0,\frac{1}{\beta})\supseteq S$. By properties (46) and (47), it suffices to show $\sigma_{S_{0}}(d)+\sigma_{B(0,\frac{1}{\beta})}(d)\geq\sigma_{S}(d)$ for every $d\in\mathcal{E}$.
Suppose $\sigma_{S}(d)$ is finite. Then, there exists a sequence $\{x_{i}\}$ with $x_{i}\in\mathrm{bdry\ }S$ such that $\lim_{i\rightarrow\infty}d^{T}x_{i}=\sigma_{S}(d)$. For each $x_{i}$, letting $\zeta_{i}\in N_{S}(x_{i})$ be some unit normal vector, the assumed smoothness ensures $B(x_{i}-\frac{1}{\beta}\zeta_{i},\frac{1}{\beta})\subseteq S$. Hence $x_{i}-\frac{1}{\beta}\zeta_{i}\in S\ominus B(0,\frac{1}{\beta})=S_{0}$. So
@@
Next, suppose $\sigma_{S}(d)=\infty$, then $S$ is unbounded. By [52, Proposition 1.4.2(a)] there exists a ray $\{z+t\Delta\mid t\geq 0\}\subseteq S$ with $d^{T}\Delta>0$. Letting $\zeta_{z}\in N_{S}(z)$ be a unit normal vector, we have $B(z-\frac{1}{\beta}\zeta_{z},\frac{1}{\beta})\subseteq S$. By convexity, one has $\cup_{t\geq 0}B(z-\frac{1}{\beta}\zeta_{z}+t\Delta,\frac{1}{\beta})\subseteq S$. Consequently, $z-\frac{1}{\beta}\zeta_{z}+t\Delta\in S_{0}$ for any $t\geq 0$. Thus
@@
A.2.4 Proof of Lemma 7
This result follows rather directly from relating the local quadratic upper and lower bounds from smoothness and strong convexity of $h$ to balls inside its epigraph. Note given any $g\in\partial h(x)$, $(\zeta,\delta)=(g,-1)/\|(g,-1)\|$ is a unit normal $(\zeta,\delta)\in N_{\mathrm{epi\ }h}(x,h(x))$. Further note $\|\zeta\|^{2}=1-\delta^{2}$. First, we address smoothness. Consider any $\tilde{L}>L$ and $\eta>0$ such that all $y\in B(x,\eta)$ have
@@ A.2.4 Proof of Lemma 7
Local $\tilde{\beta}-$smoothness of $\text{epi }h$ w.r.t. $((x,h(x)),(\zeta,\delta))$ corresponds to the following upper bound holding for all $y$ near $x$
@@
$\nabla^{2}q(x)$ $=\tilde{L}I\hskip 18.49988pt\text{and}\hskip 18.49988pt\nabla^{2}b(x)=\frac{\tilde{\beta}}{|\delta|}I+\frac{\tilde{\beta}\zeta\zeta^{T}}{|\delta|^{3}}\ .$
Thus $\nabla^{2}b(x)-\nabla^{2}q(x)$ is positive semidefinite whenever $\tilde{L}-\tilde{\beta}/|\delta|\geq 0$. So local smoothness holds with any $\tilde{\beta}\geq\tilde{L}|\delta|$-smoothness holds, and so $L\delta$ local smoothness holds.
@@
Similarly, these two Hessians can be computed as
$\nabla^{2}q(x)$ $=\tilde{\mu}I\hskip 18.49988pt\text{and}\hskip 18.49988pt\nabla^{2}b(x)=\frac{\tilde{\alpha}}{|\delta|}I+\frac{\tilde{\alpha}\zeta\zeta^{T}}{|\delta|^{3}}\ .$
The minimum eigenvalue of $\nabla^{2}q(x)-\nabla^{2}b(x)$ is $\tilde{\mu}-\tilde{\alpha}/|\delta|^{3}$ and so any $\tilde{\alpha}\leq\tilde{\mu}|\delta|^{3}$-strong convexity of $\text{epi }h$ w.r.t. $((x,h(x)),(\zeta,\delta))$ holds locally.
A.2.5 Proof of Lemma 19
Consider any $y\in\mathbb{E}$ with $\gamma_{S}(y)\neq 0$. Let $\bar{y}=y/\gamma_{S}(y)$ be the corresponding point on the boundary of $S$. Note that $\zeta\in N_{S}(\bar{y})$ if and only if the halfspace $\mathcal{H}=\{z\mid\zeta^{T}(z-\bar{y})\leq 0\}\supseteq{S}$, or equivalently $\gamma_{S}\geq\gamma_{\mathcal{H}}$. Then it yields
$N_{S}(\bar{y})$ $=\{\zeta\mid\gamma_{S}(z)\geq\gamma_{\mathcal{H}}(z),\forall z\}$
$=\{\zeta\mid\gamma_{S}(z)\geq\gamma_{S}(\bar{y})+\frac{\zeta^{T}(z-\bar{y})}{\zeta^{T}\bar{y}},\forall z\}$
$=\{\zeta\mid\gamma_{S}(z)\geq\gamma_{S}(y)+\frac{\zeta^{T}(z-y)}{\zeta^{T}\bar{y}},\forall z\}$
$=\{\zeta\mid\frac{\zeta}{\zeta^{T}\bar{y}}\in\partial\gamma_{S}(y)\}$
where the second equality is given by Example 3.1, $\gamma_{\mathcal{H}}(z)=\frac{\zeta^{T}z}{\zeta^{T}\bar{y}}=\gamma_{S}(\bar{y})+\frac{\zeta^{T}(z-\bar{y})}{\zeta^{T}\bar{y}}$, and the third equality is by rescaling.
Then by the chain rule and this gauge subgradient formula,
@@ A.2.5 Proof of Lemma 19
Define $\bar{\zeta}=r\zeta$. Then the gauge of this ball is given by
@@
$=\inf\{\lambda>0\mid(2\bar{\zeta}^{T}\bar{y}-\|\bar{y}\|^{2})\lambda^{2}+2\lambda(\bar{y}-\bar{\zeta})^{T}z-\|z\|^{2}\geq 0\}$
$=\left\{\begin{aligned} &\frac{\|z\|^{2}}{2(\bar{y}-\bar{\zeta})^{T}z},&\|\bar{y}\|^{2}=2\zeta^{T}\bar{y}\\ &\frac{-(\bar{y}-\bar{\zeta})^{T}z+\sqrt{((\bar{y}-\bar{\zeta})^{T}z)^{2}+(2\bar{\zeta}^{T}\bar{y}-\|\bar{y}\|^{2})\|z\|^{2}}}{2\bar{\zeta}^{T}\bar{y}-\|\bar{y}\|^{2}},&\|\bar{y}\|^{2}\neq 2\zeta^{T}\bar{y}\ .\end{aligned}\right$.
@@
    $\nabla^{2}\frac{1}{2}\gamma_{B}^{2}(y)=\frac{1}{(\bar{\zeta}^{T}\bar{y})^{2}}\left(3\bar{\zeta}\bar{\zeta}^{T}-(\bar{\zeta}\bar{y}^{T}+\bar{y}\bar{\zeta}^{T})+\bar{\zeta}\bar{y}I\right)\ .$
Case 2: $\|\bar{y}\|^{2}\neq 2\zeta^{T}\bar{y}$. Note
$((\bar{y}-\bar{\zeta})^{T}y)^{2}+(2\bar{\zeta}^{T}\bar{y}-\|\bar{y}\|^{2})\|y\|^{2}=(\bar{\zeta}^{T}\bar{y})^{2}\ .$ (50)
@@
    $\qquad\left.-\frac{\left(2\left((\bar{y}-\bar{\zeta})^{T}z\right)^{2}+(2\bar{\zeta}^{T}\bar{y}-\|\bar{y}\|^{2})\|z\|^{2}\right)(\bar{y}-\bar{\zeta})+(2\bar{\zeta}^{T}\bar{y}-\|\bar{y}\|^{2})(\bar{y}-\bar{\zeta})^{T}zz}{\sqrt{\left((\bar{y}-\bar{\zeta})^{T}z\right)^{2}+(2\bar{\zeta}^{T}\bar{y}-\|\bar{y}\|^{2})\|z\|^{2}}}\right)\ ,$
@@
    $\qquad-\frac{\|z\|^{2}\left((\bar{y}-\bar{\zeta})z^{T}+z(\bar{y}-\bar{\zeta})^{T}\right)}{\left(\left((\bar{y}-\bar{\zeta})^{T}z\right)^{2}+(2\bar{\zeta}^{T}\bar{y}-\|\bar{y}\|^{2})\|z\|^{2}\right)^{\frac{3}{2}}}+\frac{(\bar{y}-\bar{\zeta})^{T}zzz^{T}}{\left(\left((\bar{y}-\bar{\zeta})^{T}z\right)^{2}+(2\bar{\zeta}^{T}\bar{y}-\|\bar{y}\|^{2})\|z\|^{2}\right)^{\frac{3}{2}}}$
    $\qquad+\left(\frac{1}{2\bar{\zeta}^{T}\bar{y}-\|\bar{y}\|^{2}}-\frac{(\bar{y}-\bar{\zeta})^{T}z}{(2\bar{\zeta}^{T}\bar{y}-\|\bar{y}\|^{2})\sqrt{\left((\bar{y}-\bar{\zeta})^{T}z\right)^{2}+(2\bar{\zeta}^{T}\bar{y}-\|\bar{y}\|^{2})\|z\|^{2}}}\right)I\ .$
    At point $z=y$, the numerator in the above Hessian formula can be notably simplified by (50). After such simplifications, the Hessian matrix at $y$ is given by
@@ A.2.6 Proof of Lemma 20
$\lambda_{i}$ $=0,\hskip 18.49988pt\text{for }i=2,...,n-1$
@@
$\lambda^{2}-\left(C_{1}\|a\|^{2}+2C_{2}a^{T}b+C_{3}\|b\|^{2}\right)\lambda+\left(C_{1}C_{3}-C_{2}^{2}\right)\left(\|a\|^{2}\|b\|^{2}-(a^{T}b)^{2}\right)=0\ .$
If $a$ and $b$ are linearly dependent, then the matrix is a rank one matrix. The only nonzero eigenvalue is $C_{1}\|a\|^{2}+2C_{2}a^{T}b+C_{3}\|b\|^{2}$, which matches the formula. ∎
@@
$\lambda_{i}=$ $0+\frac{1}{\bar{\zeta}^{T}\bar{y}}=\frac{1}{\bar{\zeta}^{T}\bar{y}},\hskip 18.49988pt\text{for }i=2,...,n-1\ ,$
where $\bar{\zeta}=r\zeta$.